OpenStudy (anonymous):

find the distance from the line 4x+3y=5 to the point (6,9)

OpenStudy (anonymous):

distance equation?

OpenStudy (anonymous):

m=-4/3 y-9=-4/3(x-6) y-9=-4/3x+8 y=-4/3x +17 (0, 5/3) and (6,9) d=s.rt (6-0)^2+(9-5/3)^2 =9.47511

OpenStudy (anonymous):

perhaps try this.

OpenStudy (anonymous):

The distance from a point P(x,o,yo) to a line ax+by+c = 0 is \[d=\frac{ \left| ax _{o}+by _{o}+c \right| }{ \sqrt{a^2+b^2} }\]

OpenStudy (anonymous):

\[d=\frac{ \left| 4(6)+3(9)-5 \right| }{ \sqrt{4^2+3^2} }\]

OpenStudy (anonymous):

been a while since I've done this so please correct me if i'm wrong

OpenStudy (anonymous):

i got 9.2... but the hint say first find the line through the point (6,9) parallel to the given line, and then find the distance between these lines

OpenStudy (welshfella):

you made a slight error on the slope of the line perpendicular to the given line its slope is - 1 / (-4/3) = 3/4

OpenStudy (welshfella):

- oh the parallel line ! sorry

OpenStudy (anonymous):

so was it right or wrong?

OpenStudy (welshfella):

chrs00's method is correct and that comes to 9.2

OpenStudy (anonymous):

yep

OpenStudy (welshfella):

why did you use (0,5/3) on the original line?

OpenStudy (anonymous):

i think he just used any point on that line. i think you need to use a point that is perpendicular to that line

OpenStudy (anonymous):

distance=shortest distance

OpenStudy (anonymous):

are u sure ..this question is structure different frm the previous set of questions that i did..they want .. to find distance frm the point (a,b) eg.(4,8) to the line y=sumthing eg 6x+8y=9 ... inwhich i used that formula provide by chrs00 but the question ask to find the distance frm the line given to the point given...

OpenStudy (welshfella):

Chris00 formula gives you just that

OpenStudy (anonymous):

you were given a point that did not lie on that line

OpenStudy (welshfella):

another way to do it is to find the equation of the line perpendicular to the original line and passing through the point (6,9) then solving the 2 equations simultaneously to find the point of intersection. Then use the distance formula with this point and (6,9) But that's a bit long-winded.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u sketch it?

OpenStudy (welshfella):

you can sketch it on the desmos web site

OpenStudy (anonymous):

but i wouldnt get desmos to use in my test

OpenStudy (welshfella):

the 2 parallel lines will look something like 

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

the line joining the point (0, 5/3) and (6,9) is not perpendicular to the 2 parallel lines. That is why you didn't get the right answer